The probability is 0.4 that a traffic fatality involves an intoxicated or​ alcohol-impaired driver or nonoccupant. In eight traffic​ fatalities, find the probability that the​ number, Y, which involve an intoxicated or​ alcohol-impaired driver or nonoccupant is

d. Obtain the standard deviation of Y.
Answer #1

Answer

Sample size is n = 8

probability = 0.4 = p

We know that the formula for the standard deviation is given as

Standard deviation =  \sqrt{p*(1-p)*n}

setting the given values, we get

Standard deviation = \sqrt{0.4*(1-0.4)*8} = \sqrt{1.92} = 1.39

So, the required standard deviation is 1.39 (rounded to two decimals)